Sarah lives where weekly unemployment comp. is 55% of the 26-week avg. for the two highest-salaried quarters. If Sarah made an annual salary of $34,000 before being laid off, what will be her weekly unemployment compensation?


1
Expert's answer
2021-09-24T13:58:50-0400

Assuming annual salary before being laid-off was earned evenly through-out the year.

Annual salary = $34,000

Salary for 26 weeks or 2 quarters "= ( \frac{\$34000 }{ 4} ) \times 2 = 17000"

Average salary (per week)"= \frac{17000 }{26 \space weeks} = \$653.846" per week

Therefore, weekly un-employement compensation"= 55\% \times 653.846 = \$359.615"